OpenStudy (anonymous):

type this equation in standard form.3=2y+7x @Bea_26 @Briaja_Summers

OpenStudy (uri):

-7x - 2y = -3 Signs will change and all the values which are in left will come to right and vise versa.

OpenStudy (anonymous):

3=2y+7x is y=x+1. Okay this is really easy, dont you have like a book or website to like get this from.? You should learn first, then if you still dont get it you can ask me or whoever.

OpenStudy (anonymous):

First i dont know if i did this right though, so haha stop asking i might get them wrong.. but i tried.